Ethanol has a normal boiling point of 78 °C and a heat of vaporization of 38.56 kJ/mol. At what pressure (in atm) would it have a boiling point of 85.0 °C?


Determine the amount of heat required to convert 500.0 g of liquid ethanol at 25.0°C into gaseous ethanol at 125°C. Use the following information on ethanol (CHOH) to calculate the amount of heat for each of the steps.

Melting Point = −117°C Boiling Point = 78.3°C

Molar Heat Capacities:

C =113 J/mol °C

C = 420 J/mol °C

ΔH = 5.02 kJ/mol

ΔH = 39.3 kJ/mol


Which of these transitions are exothermic processes? Multiple answers are accepted.

a. Na(g) → Na(l)

b. Na(s) → Na(g)

c. Na(g) → Na(g) + e

d. Cl(g) → 2Cl(g)

e. Cl(g) + e→ Cl(g)



1
Expert's answer
2021-04-08T12:19:50-0400

"T_1=(273+38.56)=311.56" "P_1=?"

"T_2=(273+78)=351k" "P_2=760Torr"


"ln(\frac{P_2}{P_1})=\frac{∆vap}{R}[\frac{1}{T_1}-\frac{1}{T_2}]"


"ln(\frac{760}{P_1})=\frac{38560}{8.314}[\frac{1}{311.56}-\frac{1}{351}]"


"ln(\frac{760}{P_1})=4638×3.60×10^{-4}= 1.67" =


"ln(\frac{760}{P_1})=e^{1.67}=5.312"


"P_1=\frac{760}{5.312}=143.07" Torr






q = m∆Hvap = 32.6 g x 879 J/g = 28,655 J






a. Na(g) → Na(l)

b. Na(s) → Na(g)

c. Na(g) → Na(g) + e